OpenStudy (anonymous):

Find the power expansion for integral 1/(x^7 + 1) dx for |x| < 1.

OpenStudy (anonymous):

\[\frac{1}{1-x}=\sum_{n=0}^{\infty}x^n\] replace \(x\) by \(-x^7\)

OpenStudy (anonymous):

first step is to write \[\sum_{n=0}^{\infty}((-x)^7)^n\] the clean up with some algebra

OpenStudy (anonymous):

could write it as \[\sum_{n=0}^{\infty}(-1)^nx^{7n}\]

OpenStudy (anonymous):

oop it was the integral. so there is one more step, take the expansion written above and integrate it term by term

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw